# DeepAgent: A General Reasoning Agent with Scalable Toolsets

### v1

@@ 4.3. Implementation Details
We use QwQ-32B (Team, 2024) as DeepAgent’s backbone model, with Qwen2.5-32B-Instruct (Qwen et al., 2024) as the auxiliary model in our main results. Text generation employs a maximum of 81,920 tokens with temperature 0.7, top_p 0.8, top_k 20, and repetition penalty 1.05. Web search and page browsing are implemented using Google Serper API and Jina Reader API, respectively. The VQA tool is based on Qwen2.5-VL-32B-Instruct (Bai et al., 2025). Tool retrieval is performed using bge-large-en-v1.5 (Xiao et al., 2024). Training consists of 100 steps of ToolPO with batch size 64, $\lambda_{1}=\lambda_{2}=1$, rollout size $K=8$, and maximum sequence length 32,768. Additional details are provided in Appendix C. All experiments are conducted on 64 NVIDIA H20-141GB GPUs.
@@ 5.5. Effectiveness of Tool Retrieval Strategies
To compare pre-retrieving tools versus autonomous discovery during task execution, we conduct experiments shown in Table 4. (1) The on-demand nature of dynamic tool discovery yields superior performance and robust scalability. Autonomous tool retrieval during reasoning consistently outperforms pre-retrieved tools across all frameworks, demonstrating the superiority of on-demand tool access in open-set scenarios. Performance gains are most pronounced on large toolsets like ToolBench (16k tools) and ToolHop (3.9k tools), indicating robust scalability for real-world tasks. (2) DeepAgent synergizes better with dynamic retrieval. Combined with autonomous tool retrieval, our framework achieves the best results by a large margin, scoring 52.6 on average versus 28.5 for the best workflow-based method. This demonstrates DeepAgent’s architecture is uniquely suited for dynamic tool discovery.
@@ 5.6. Scaling Analysis of Action Limits
Table 5. Performance comparison with different reasoning model backbones, spanning MOE-based models with 30B and 235B parameters.
@@ 6. Conclusion
In this work, we introduce DeepAgent, an end-to-end reasoning agent that unifies thinking, tool discovery, and execution into a single, coherent agentic reasoning process. To enable robust long-horizon interaction, we propose an autonomous memory folding mechanism that compresses interaction history into a structured memory, allowing the agent to ”take a breath” and reconsider its strategy. We also introduce ToolPO, an end-to-end RL method that leverages LLM simulated APIs for stable training and fine-grained advantage attribution for precise credit assignment to tool invocations. Extensive experiments on general tool-use and downstream applications demonstrate that DeepAgent significantly outperforms various baseline agents, particularly in open-set scenarios requiring dynamic tool discovery over scalable toolsets.
@@ A.2. Benchmarks
  • ToolHop (Ye et al., 2025a): A multi-hop reasoning dataset comprising 995 complex questions. It leverages 3,912 locally executable tools and requires between 3 to 7 sequential tool calls per task.
@@
Appendix B Baselines
We compare our proposed method with several baseline agents. The details of these baselines are introduced as follows:
  • ReAct (Reasoning and Acting) (Yao et al., 2022b): ReAct is a general paradigm that combines reasoning and acting with language models. It prompts the model to generate a sequence of interleaved thought, action, and observation steps to solve a given task.
  • CodeAct (Wang et al., 2024a): This is a framework where the agent’s actions are expressed as Python code, which are then executed in an interpreter. By using code as the action space, the agent can interact with a wide variety of tools, APIs, and system functionalities.
  • Plan-and-Solve (Wang et al., 2023): This method follows a two-stage process to tackle complex problems. First, the model devises a detailed, step-by-step plan to solve the problem without using any tools. Then, it executes the plan, carrying out the necessary calculations or actions as outlined.
  • Reflexion (Shinn et al., 2023): Reflexion is an approach that enhances agent learning through verbal self-reflection. After a failed attempt, the agent reflects on what went wrong and records this reflection in its memory.
  • AgentLM (Zeng et al., 2024): An instruction tuning method designed to enhance the general agent capabilities of LLMs. It uses a lightweight, specially curated dataset called AgentInstruct to fine-tune LLMs.
  • WebThinker (Li et al., 2025b): WebThinker is a deep research agent designed for complex information-seeking tasks. It autonomously explores the web by interleaving a ”think-search-write” process to gather and synthesize information.
  • HiRA (Hierarchical Reasoning Agent) (Jin et al., 2025a): HiRA employs a hierarchical agent architecture to solve complex, multi-modal tasks. It decouples high-level planning from low-level execution by using a planner to decompose the task and executor agents to perform specific actions like searching or coding.
  • OpenAI Deep Research (OpenAI, 2025): This is a feature within ChatGPT that performs in-depth research on complex topics by exploring a wide range of web sources. It takes more time than a standard query to synthesize information and generate a detailed, structured response.
Appendix C Implementation Details
@@
Appendix D Memory Schema
Our brain-inspired memory architecture consists of three components: episodic, working, and tool memory. To ensure stable memory folding and prevent information loss, each component is defined by a specific JSON schema. This structured format enables the agent to reliably parse and utilize the compressed memory, facilitating robust long-term reasoning.
@@ Appendix D Memory Schema
Episodic memory provides a high-level summary of the agent’s task progression, major milestones, decisions, and outcomes. This allows the agent to maintain long-term context and reflect on its overall strategy. The format is: {”task_description”: ”A general summary of what the reasoning history has been doing and the overall goals it has been striving for.”, ”key_events”: [{”step”: ”step number”, ”description”: ”A detailed description of the specific action taken, decision made, or milestone achieved at this step, including relevant context and reasoning behind the choice.”, ”outcome”: ”A detailed account of the direct result, observation, or feedback received from this action or decision, including any new information gained or changes in the task state.”}], ”current_progress”: ”A general summary of the current progress of the task, including what has been completed and what is left to be done.”}
@@
Working memory functions as the agent’s short-term buffer, holding information relevant to its immediate context. It focuses on the current sub-goal, active challenges, and planned next steps, ensuring continuity of reasoning across memory folds. The format is: {”immediate_goal”: ”A clear summary of the current subgoal—what you are actively working toward at this moment.”, ”current_challenges”: ”A concise summary of the main obstacles or difficulties you are presently encountering.”, ”next_actions”: [{ẗype”: ”tool_call or planning ordecision”, d̈escription”: ”Anticipate and describe the next concrete action you intend to take to advance the task.”}]}
@@
Tool memory consolidates the agent’s experiences with various tools. It tracks usage patterns, success rates, effective parameter combinations, and common errors. This structured knowledge enables the agent to learn from its interactions and refine its tool-use strategies over time. The format is: {”tools_used”: [{”tool_name”: ”string”, ”success_rate”: ”float”, ”effective_parameters”: [”param1”, ”param2”], ”common_errors”: [”error_type1”, ”error_type2”], ”response_pattern”: ”description of typical output”, ”experience”: ”Reflect and summarize your experience, including both successes and failures.”}], ”derived_rules”: [”When X condition occurs, prefer tool Y”, ”Tool Z works best with parameter A set to B”]}
Appendix E Case Study

### v2

@@ 4.3. Implementation Details
We use QwQ-32B (Team, 2024) as DeepAgent’s backbone model, with Qwen2.5-32B-Instruct (Qwen et al., 2024) as the auxiliary model in our main results. Text generation employs a maximum of 81,920 tokens with temperature 0.7, top_p 0.8, top_k 20, and repetition penalty 1.05. Web search and page browsing are implemented using Google Serper API and Jina Reader API, respectively. The VQA tool is based on Qwen2.5-VL-32B-Instruct (Bai et al., 2025). Tool retrieval is performed using bge-large-en-v1.5 (Xiao et al., 2024). Training consists of 100 steps of ToolPO with batch size 64, $\lambda_{1}=\lambda_{2}=1$, rollout size $K=8$, and maximum sequence length 32,768. Additional details are provided in Appendix B. All experiments are conducted on 64 NVIDIA H20-141GB GPUs.
@@ 5.5. Effectiveness of Tool Retrieval Strategies
To compare pre-retrieving tools versus autonomous discovery during task execution, we conduct experiments shown in Table 4. We find: (1) The on-demand nature of dynamic tool discovery yields superior performance and robust scalability. Autonomous tool retrieval during reasoning consistently outperforms pre-retrieved tools across all frameworks, demonstrating the superiority of on-demand tool access in open-set scenarios. Performance gains are most pronounced on large toolsets like ToolBench (16k tools) and ToolHop (3.9k tools), indicating robust scalability for real-world tasks. (2) DeepAgent synergizes better with dynamic retrieval. Combined with autonomous tool retrieval, our framework achieves the best results by a large margin, scoring 52.6 on average versus 28.5 for the best workflow-based method. This demonstrates that DeepAgent’s architecture is uniquely suited for dynamic tool discovery.
@@ 5.6. Scaling Analysis of Action Limits
Table 5. Performance with different reasoning model backbones: MOE-based models with 30B and 235B parameters.
@@ 6. Conclusion
In this work, we introduce DeepAgent, an end-to-end reasoning agent that unifies thinking, tool discovery, and execution into a single, coherent agentic reasoning process. To enable robust long-horizon interaction, we propose an autonomous memory folding mechanism that compresses interaction history into a structured memory, allowing the agent to ”take a breath” and reconsider its strategy. We also introduce ToolPO, an end-to-end RL method that leverages LLM simulated APIs for stable training and fine-grained advantage attribution for precise credit assignment to tool invocations. Extensive experiments on general tool-use and downstream applications demonstrate that DeepAgent significantly outperforms various baseline agents, particularly in open-set scenarios requiring dynamic tool discovery over scalable toolsets. This work opens new avenues for creating more general and capable agents, with future research directions including developing advanced memory management, multimodal (or even omni-modal) capabilities, and human-agent collaborative reasoning.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China No. 62272467, and the China Postdoctoral Science Foundation under Grant Number 2025T180440. The work was partially done at the Engineering Research Center of Next-Generation Intelligent Search and Recommendation, MOE.
@@ A.2. Benchmarks
  • ToolHop (Ye et al., 2025): A multi-hop reasoning dataset comprising 995 complex questions. It leverages 3,912 locally executable tools and requires between 3 to 7 sequential tool calls per task.
@@
Appendix B Implementation Details
@@
Appendix C Memory Schema
Our brain-inspired memory architecture contains three components: episodic, working, and tool memory. To support stable memory folding and reduce information loss, we define each component with a fixed JSON schema, enabling reliable parsing and use of compressed memories for long-horizon reasoning.
@@ Appendix C Memory Schema
Episodic memory records high-level task progression (milestones, decisions, outcomes) to preserve long-term context. The format is: {”task_description”: ”A general summary of what the reasoning history has been doing and the overall goals it has been striving for.”, ”key_events”: [{”step”: ”step number”, ”description”: ”A detailed description of the specific action taken, decision made, or milestone achieved at this step, including relevant context and reasoning behind the choice.”, ”outcome”: ”A detailed account of the direct result, observation, or feedback received from this action or decision, including any new information gained or changes in the task state.”}], ”current_progress”: ”A general summary of the current progress of the task, including what has been completed and what is left to be done.”}
@@
Working memory captures the immediate goal, active challenges, and next actions to maintain continuity across folds. The format is: {”immediate_goal”: ”A clear summary of the current subgoal—what you are actively working toward at this moment.”, ”current_challenges”: ”A concise summary of the main obstacles or difficulties you are presently encountering.”, ”next_actions”: [{ẗype”: ”tool_call or planning ordecision”, d̈escription”: ”Anticipate and describe the next concrete action you intend to take to advance the task.”}]}
@@
Tool memory consolidates tool-use experience (success rates, effective parameters, common errors) and derives reusable rules. The format is: {”tools_used”: [{”tool_name”: ”string”, ”success_rate”: ”float”, ”effective_parameters”: [”param1”, ”param2”], ”common_errors”: [”error_type1”, ”error_type2”], ”response_pattern”: ”description of typical output”, ”experience”: ”Reflect and summarize your experience, including both successes and failures.”}], ”derived_rules”: [”When X condition occurs, prefer tool Y”, ”Tool Z works best with parameter A set to B”]}
Appendix D Case Study